If 7.46 mg of a gas occupies 2899 mL at 849 mm Hg and 27.63°C, what is the molecular weight of the gas?



1
Expert's answer
2021-12-06T14:20:02-0500


V= 2.899litres

P = 849 mmHg

T=273+27.63=300.63k

: R = 8.314 J/mol·K.


PV=nRT


n"=\frac{849×2.899}{300.63×8.314}=0.98"


Molecular weight =